Given the bulk modulus of sea water is 2.2*10^9Pa




and the density of



surface sea water is 1.025*10^3 kg/m^3




Calculate the volume of



1.025*10^3 kg of sea water at a depth near



5 km where the pressure is



5.0*10^7 Pa.

1
Expert's answer
2022-01-21T13:49:55-0500
"\dfrac{ΔV}V=\dfrac{ΔP}B,\\\space\\

\dfrac{V-V'}V=\dfrac{ΔP}{B},\\\space\\

V'=V\bigg(1-\dfrac{ΔP}B\bigg)=\dfrac mρ ·\bigg(1-\dfrac{ΔP}B\bigg),\\\space\\

V'=(1.025·10^3)/(1.025·10^3 ) ·(1-(5.0·10^7)/(2.2·10^9 ))=0.98\text{ m}^3"